# A Question-centric Multi-experts Contrastive Learning Framework for Improving the Accuracy and Interpretability of Deep Sequential Knowledge Tracing Models

### v1

@@ 5.5. Ablation Study (RQ2)
Table 3. The AUC and Accuracy performance of different variants in our proposed Q-MCKT. Marker ***, $\circ$ and $\bullet$ indicates whether our proposed model is statistically superior/equal/inferior to the compared method (using paired t-test at 0.01 significance level).
@@ 5.6. Analysis of Contrastive Learning (RQ3)
Ours 0.7789±0.0019 0.7805±0.0007 0.7098±0.0024 0.7354±0.0021 0.7213±0.0004 0.7032±0.0012
w/o CL 0.7765±0.0016* 0.7716±0.0005* 0.7027±0.0015* 0.7345±0.0015* 0.7175±0.0006* 0.7008±0.0015*
Table 4. The AUC and Accuracy performance of Q-MCKT on less interactive questions. Marker ***, $\circ$, and $\bullet$ indicate whether our proposed model is statistically superior/equal/inferior to the compared method (using paired t-test at 0.01 significance level).
@@
To further investigate the effectiveness of our proposed contrastive learning procedure, we conduct additional experiments to analyze the impact of contrastive learning specifically on less interactive questions [^17] To save computation, we only experiment on the three datasets, as the performance with and without CL in Algebra2005 dataset is not significantly different.. As presented in Table 4, the results demonstrate that our proposed contrastive learning procedure helps the less interactive question learn a more general representation under relative rare updated times. This more general representation leads to more robust and accurate question knowledge acquisition state for the prediction layer to get a better performance (e.g., an improvement in AUC of 0.89% in NIPS34).
@@ 5.7. Impact of Expert Number (RQ4)
1 0.7822±0.0017* 0.8361±0.0008* 0.8014±0.0002* 0.7211±0.0020* 0.7357±0.0013* 0.8222±0.0006* 0.7321±0.0005* 0.7059±0.0013*
2 0.7864±0.0014* 0.8429±0.0011 0.8035±0.0004 0.7221±0.0022* 0.7365±0.0009* 0.8239±0.0006 0.7329±0.0005 0.7058±0.0015*
3 0.7876±0.0016 0.8418±0.0005* 0.8026±0.0003* 0.7222±0.0025* 0.7376±0.0011 0.8229±0.0006* 0.7323±0.0004* 0.7061±0.0013*
4 0.7873±0.0014* 0.8421±0.0008* 0.8027±0.0004* 0.7228±0.0023 0.7370±0.0007* 0.8221±0.0006* 0.7323±0.0005* 0.7067±0.0008
5 0.7874±0.0017* 0.8419±0.0006* 0.8026±0.0004* 0.7223±0.0025* 0.7366±0.0010* 0.8225±0.0005* 0.7319±0.0002* 0.7058±0.0014*
Table 5. The AUC and Accuracy performance of different numbers of experts in Q-MCKT. Marker ***, $\circ$, and $\bullet$ indicate whether our proposed model is statistically superior/equal/inferior to the compared method (using paired t-test at 0.01 significance level).
In this section, we investigate the impact of the number of experts on the performance of knowledge tracing (KT). The variable $e$ represents the number of experts, and we conduct experiments ranging from 1 to 5 experts. As shown in Table 5, we observe that as the number of experts increases, the model’s performance improves slightly (e.g., from $e=2$ to $e=4$ in the EdNet dataset) or even experiences a slight degradation (e.g., from $e=2$ to $e=5$ in the Algebra2005 and NIPS34 datasets). This observation suggests that the optimal number of experts may vary for different datasets. Moreover, the overall differences in performance among the cases with numbers of experts greater than 2 are not deemed significant. We attribute this phenomenon to the limited size of the training data in our experimental datasets, preventing the effective learning of all the experts (See Table 1). Specifically, in the MoE technique, the gated mechanism aggregates each expert’s output to derive an overall output. In this context, despite an increase in the number of experts, the overall output is mainly influenced by only a few of them.
@@ 5.8. Visualization of Knowledge State Estimation (RQ5)
Figure 8. An example of a student knowledge states of 5 concepts as student solves 50 questions of Algebra2005.
To verify the accurate estimations of students’ knowledge states by the Q-MCKT model, we randomly select a student sequence from Algebra2005181818As its Avg. per Q is almost equal to 1. and observe the knowledge state variations of the student in 50 questions with 5 KCs. From Figure 8, we observe that: (1) the estimated knowledge acquisition state of $c_{2}$ decreases quite a bit when the student mistakenly answers the questions that contain the KC $c_{2}$ twice (e.g. questions 5 and 8). (2) On the other hand, the student gives the right answers to the questions (questions 22, 26, and 31) which are related to the KC $c_{3}$ hence the knowledge acquisition of KC $c_{3}$ is constantly increasing. (3) As the student practices more questions, the knowledge acquisition state estimations become much more stable and after finishing all these 50 questions, the model is confident that the student has acquired the KCs $c_{1}$, $c_{3}$ and $c_{5}$.

### v2

@@ 5.5. Ablation Study (RQ2)
Table 3. The AUC and Accuracy performance of different variants in our proposed Q-MCKT. We highlight the highest results with bold. Marker *, $\circ$ and $\bullet$ indicates whether our Q-MCKT model is statistically superior/equal/inferior to its different variants (using paired t-test at 0.01 significance level).
@@ 5.6. Analysis of Contrastive Learning (RQ3)
Table 4. The AUC and Accuracy performance of Q-MCKT and its w/o CL variant on less interactive questions. We highlight the highest results with bold. Marker *, $\circ$, and $\bullet$ indicate whether our proposed model is statistically superior/equal/inferior to the variant (using paired t-test at 0.01 significance level).
@@
Q-MCKT 0.7789±0.0019 0.7805±0.0007 0.7098±0.0024 0.7354±0.0021 0.7213±0.0004 0.7032±0.0012
w/o CL 0.7765±0.0016* 0.7716±0.0005* 0.7027±0.0015* 0.7339±0.0017* 0.7158±0.0005* 0.7001±0.0014*
@@
To further investigate the effectiveness of our proposed contrastive learning procedure, we conduct additional experiments to analyze the impact of contrastive learning specifically on less interactive questions [^12] To save computation, we only experiment on the three datasets, as the performance with and without CL in Algebra2005 dataset is comparatively less significant when compared to the other three datasets.. As presented in Table 4, the results demonstrate that our proposed contrastive learning procedure helps the less interactive question learn a more general representation under relative rare updated times. This more general representation leads to more robust and accurate question knowledge acquisition state for the prediction layer to get a better performance (e.g., an improvement in AUC of 0.89% in NIPS34).
@@ 5.7. Impact of Expert Number (RQ4)
Table 5. The AUC and Accuracy performance of different numbers of experts in Q-MCKT. We highlight the highest results with bold.
@@
1 0.7822±0.0017 0.8361±0.0008 0.8014±0.0002 0.7211±0.0020 0.7332±0.0014 0.8204±0.0006 0.7314±0.0005 0.7049±0.0014
2 0.7864±0.0014 0.8429±0.0011 0.8035±0.0004 0.7221±0.0022 0.7365±0.0009 0.8239±0.0006 0.7329±0.0005 0.7058±0.0015
3 0.7876±0.0016 0.8418±0.0005 0.8026±0.0003 0.7222±0.0025 0.7376±0.0011 0.8229±0.0006 0.7323±0.0004 0.7061±0.0013
4 0.7873±0.0014 0.8421±0.0008 0.8027±0.0004 0.7228±0.0023 0.7373±0.0007 0.8226±0.0006 0.7323±0.0005 0.7067±0.0008
5 0.7874±0.0017 0.8419±0.0006 0.8026±0.0004 0.7223±0.0025 0.7371±0.0010 0.8225±0.0005 0.7321±0.0002 0.7062±0.0014
In this section, we investigate the impact of the number of experts on the performance of KT. The variable $e$ represents the number of experts, and we conduct experiments ranging from 1 to 5 experts. As shown in Table 5, we observe that as the number of experts increases, the performance of the model improves accordingly (e.g., from $e=1$ to $e=3$ in the ASSIST2009 dataset). Moreover, it is noteworthy that once the performance reaches its peak, the inclusion of additional experts does not significantly affect the performance. This observation can be attributed to the fact that, when the training data size is fixed, there exists an associated optimal number of learnable experts that effectively capture the important patterns, while the additional experts may not acquire new information and therefore have minimal impact on the overall output. During the inference stage, in the MoE technique, the gated mechanism aggregates the outputs of all the experts to derive an overall output. However, despite the inclusion of additional experts, the overall output is primarily influenced by a subset of experts that exhibit strong contributions.
@@ 5.8. Visualization of Knowledge State Estimation (RQ5)
Figure 8. The visualization of students’ knowledge states in 50 questions with 5 concepts.
To verify the accurate estimations of students’ knowledge states by the Q-MCKT model, we randomly select a student sequence from Algebra2005131313As its Avg. per Q is almost equal to 1. and observe the knowledge state variations of the student in 50 questions with 5 KCs. From Figure 8, we observe that: (1) the estimated knowledge acquisition state of $c_{2}$ decreases quite a bit when the student mistakenly answers the questions that contain the KC $c_{2}$ twice (e.g. questions 5 and 8). (2) On the other hand, the student gives the right answers to the questions (questions 22, 26, and 31) which are related to the KC $c_{3}$ hence the knowledge acquisition of KC $c_{3}$ is constantly increasing. (3) As the student practices more questions, the knowledge acquisition state estimations become much more stable and after finishing all these 50 questions, the model is confident that the student has acquired the KCs $c_{1}$, $c_{3}$ and $c_{5}$.
5.9. Visualization of Prediction Results (RQ6)
Figure 9. The visualization of Q-MCKT’s prediction results in 50 questions.
We also randomly select a student sequence from ASSIST2009 and conduct an additional visualization to demonstrate how our Q-MCKT model produces the interpretable prediction results. Figure 9 (a) illustrates the direct prediction result, denoted as $\hat{r}$. Figure 9 (b) shows the corresponding interpretable components of the prediction result $\hat{r}$, i.e., question knowledge acquisition score $\sigma(\alpha)$ and concept knowledge acquisition score $\sigma(\bar{\beta})$, generated by the Q-MCKT model. In previous approaches, prediction results $\hat{r}$ were directly derived from the hidden states of neural networks. In contrast, our Q-MCKT model generates prediction results based on two interpretable components, which helps to elucidate the underlying rationale behind the prediction results. For instance, as shown in Figure 9 (a), the model predicts a low value of $\hat{r}$ on question 25. By examining the corresponding interpretable components in Figure 9 (b), we observe that the student exhibits a low value of question knowledge acquisition score $\sigma(\alpha)$ and a moderate value of concept knowledge acquisition score $\sigma(\bar{\beta})$ on question 25. This indicates that the student only has a moderate mastery of the KCs related to question 25, and the question’s difficulty level is relatively high for the student acquisition. Consequently, the model estimates that the student is likely to answer question 25 incorrectly.
@@ 6. Conclusion
Acknowledgements.
This work was supported in part by National Key R&D Program of China, under Grant No. 2022YFC3303600 and in part by Key Laboratory of Smart Education of Guangdong Higher Education Institutes, Jinan University (2022LSYS003).